OpenStudy (anonymous):

log5 (x+2)- log5 10=log5 100 how do you solve this

OpenStudy (anonymous):

(log(x+2))/(log(5))-(log(10))/(log(5)) = (log(100))/(log(5)) Add (log(10))/(log(5)) to both sides: (log(x+2))/(log(5)) = (log(10))/(log(5))+(log(100))/(log(5)) Divide both sides by 1/(log(5)): log(x+2) = log(5) ((log(10))/(log(5))+(log(100))/(log(5))) Cancel logarithms by taking exp of both sides: x+2 = 1000 Subtract 2 from both sides: x = 998 Now test that this solution is appropriate by substitution into the original equation: Check the solution x = 998: (log(x+2))/(log(5))-(log(10))/(log(5)) => -(log(10))/(log(5))+(log(2+998))/(log(5)) = (2 log(2)+2 log(5))/(log(5)) ~~ 2.86135 (log(100))/(log(5)) => (log(100))/(log(5)) = (log(100))/(log(5)) ~~ 2.86135 So the solution is correct. Thus, the solution is: x = 998

OpenStudy (anonymous):

log(ab)=loga + logb using this formula for your problem : log5+ log(x+2)-log5-log10=log5+log100 now individually put values of log5, log10 , log 100 , and then take anti log of x+2 There might be some other easier way ,but as per now , I could recall this much only.

OpenStudy (anonymous):

Solution posted by @DHASHNI is absolutely correct and shortcut too. Go for it ! :-)